# DPMM-CFL: Clustered Federated Learning via Dirichlet Process Mixture Model nonparametric clustering

### v1

@@ 1 Introduction
In this work, we propose to remove the need for fixing $K$ by incorporating a Bayesian nonparametric framework into CFL. Specifically, we consider a Dirichlet Process prior over cluster parameters, allowing the number of clusters to be inferred directly from data. This provides a principled way to adapt $K$ across federated rounds while optimizing per-cluster FL objectives. Our contribution is thus a CFL formulation with Bayesian inference over cluster structure, combined with standard federated optimization for learning cluster models. Section 2 presents the proposed CFL algorithm and how it is tightly coupled with the Bayesian clustering solution. Section 3 provides theoretical and implementation details of the clustering method and Section 4 discusses experimental results validating the proposed method. The article concludes with final remarks in Section 5.
@@ 2 Clustered federated learning with Bayesian clustering
  • Aggregation and distribution: for each cluster $k\leq K_{t}$, the server aggregates the model updates from corresponding clients, $\bm{\Omega}_{k}^{t}=\frac{1}{m_{k}^{t}}\sum_{i:\,c_{i}^{\,t}=k}\bm{\omega}_{i}^{t}$, and redistributes $\bm{\Omega}_{k}^{t}$ among those clients.
@@
2:  Initialize: set $K_{0}=1$, random cluster model $\bm{\Omega}_{1}^{0}$, assignments $\{c_{i}^{0}=1\}_{i=1}^{M}$, and number of clients per cluster $m_{1}^{0}=M$
@@
12:      Set $m_{k}^{t}\leftarrow\sum_{i=1}^{M}\mathbb{I}\{c_{i}^{\,t}=k\}$
13:      Update cluster model ((3))
@@ 4.2 System Settings
For the Bayesian nonparametric model, the DP concentration parameter is set to $\alpha=1.0$, and the base measure $G_{0}$ is a spherical Gaussian $\mathcal{N}(\bm{\mu}_{0},{\sigma}_{0}\mathbf{I})$ with $\bm{\mu}_{0}=\bm{0}$ and ${\sigma}_{0}^{2}=1$. Results are reported in terms of micro-accuracy ($\%$) and macro-F1, both averaged across clients. Each experiment is repeated over five Monte Carlo runs with different random seeds. Mean values are reported from the last three rounds, and shaded regions in Fig. 1 indicate one standard deviation.

### v2

@@ 1 Introduction
In this work, we remove the need to fix $K$ by incorporating a Bayesian nonparametric framework into CFL. Specifically, we place a Dirichlet Process prior over cluster parameters, so the effective number of clusters and client assignments are inferred directly from data. This yields a threshold-free clustering mechanism driven by posterior inference, allowing clusters to split and merge adaptively during training. Our contribution is thus a CFL formulation that couples Bayesian inference over cluster structure with standard federated optimization to learn one model per inferred cluster. Section 2 presents the proposed CFL algorithm and how it is tightly coupled with the Bayesian clustering solution. Section 3 provides theoretical and implementation details of the clustering method and Section 4 discusses experimental results validating the proposed method. The article concludes with final remarks in Section 5.
@@ 2 Clustered federated learning with Bayesian clustering
  • Aggregation and distribution: for each cluster $k\leq K_{t}$, the server aggregates the model updates from corresponding clients, $\bm{\Omega}_{k}^{t}=\sum_{i:\,c_{i}^{\,t}=k}\frac{n_{i}}{N_{k}^{t}}\bm{\omega}_{i}^{t}$, where $N_{k}^{t}=\sum_{i:\,c_{i}^{\,t}=k}n_{i}$, and redistributes $\bm{\Omega}_{k}^{t}$ among those clients.
@@
2:  Initialize: set $K_{0}=1$, random cluster model $\bm{\Omega}_{1}^{0}$ and assignments $\{c_{i}^{0}=1\}_{i=1}^{M}$
@@
12:      Set $N_{k}^{t}\leftarrow\sum_{i:\,c_{i}^{\,t}=k}n_{i}$
13:      Update cluster model ((3)): $\bm{\Omega}_{k}^{t}\leftarrow\sum_{i:\,c_{i}^{\,t}=k}\frac{n_{i}}{N_{k}^{t}}\bm{\omega}_{i}^{t}$
@@ 4.2 System Settings
For the Bayesian nonparametric model, the DP concentration parameter is set to $\alpha=1.0$ (weakly informative), and the base measure $G_{0}$ is a spherical Gaussian $\mathcal{N}(\bm{\mu}_{0},{\sigma}_{0}\mathbf{I})$ with $\bm{\mu}_{0}=\bm{0}$ and ${\sigma}_{0}^{2}=1$. Results are reported in terms of micro-accuracy ($\%$) and macro-F1, both averaged across clients. Each experiment is repeated over five Monte Carlo runs with different random seeds. Mean values are reported from the last three rounds, and shaded regions in Fig. 1 indicate one standard deviation.